OpenStudy (anonymous):

hey guys I have intense need of help with two pooptastic word problems. Anyone feel up to it??????hu huhuhu? I know you want to solve it.... The dimensions of a rectangle are such that its L is 5 inches more that its W. If the L were doubled and the W were decreased by 2, the area would be increased by 162 inches. What are the L and W of the rectangle?

OpenStudy (anonymous):

this is actaully a repeat from the other day. is this an on line class?

OpenStudy (anonymous):

Is it?

OpenStudy (anonymous):

where is the first repeat lol

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

WOW?

OpenStudy (anonymous):

can someone clue me in I feel like im the odd man ?woman out:(

OpenStudy (lgbasallote):

L = 5 +W (2L)(W-2) = 162 2L =162/w-2 L = 81/w-2 5+ W = 81/w-2 5w - 10 + w^2 - 2w = 81 w^2 + 3w -10 = 81 (w+5)(w-2) = 81 w = 76 w = 84 L = 81 when width is 76 L = 89 when width is 81 am i right?

OpenStudy (anonymous):

in any case since L is 5 inches more than W, we can write \[L=W+5\] and then double L to get \[2(W+5)=2W+10\] decrease W by 2 to get \[W-2\] and then write \[W(W+5) \] as the first area, \[(W-2)(2W+10)\] as the second area and we have \[W(W+5)+162=(W-2)(2W+10)\] and solve for W

OpenStudy (anonymous):

You guys make me fel so inadequate....thankzxxx

OpenStudy (anonymous):

ii get \[W=13\] as an answer hmm

OpenStudy (anonymous):

maybe i did something wrong, let me see

OpenStudy (anonymous):

if i write it with paper and pencil maybe i won't mess up

OpenStudy (mertsj):